Assume that 83% of the girls in a town attend college regularly. What is the standard deviation of the sampling distribution of the proportion of the girls attending college regularly if the sample size is 144? Enter your answer as a decimal, rounded to three decimal places.
Answer #1

Solution

Given that,

p = 0.83

1 - p = 1 - 0.83 = 0.17

n = 144

= p = 0.83

=  [p( 1 - p ) / n] = [(0.83 * 0.17) / 144 ] = 0.031